# A Comprehensive Framework to Operationalize Social Stereotypes for Responsible AI Evaluations

### v2

@@ Abstract
Societal stereotypes are at the center of a myriad of responsible AI interventions targeted at reducing the generation and propagation of potentially harmful outcomes. While these efforts are much needed, they tend to be fragmented and often address different parts of the issue without taking in a unified or holistic approach about social stereotypes and how they impact various parts of the machine learning pipeline. As a result, it fails to capitalize on the underlying mechanisms that are common across different types of stereotypes, and to anchor on particular aspects that are relevant in certain cases. In this paper, we draw on social psychological research, and build on NLP data and methods, to propose a unified framework to operationalize stereotypes in generative AI evaluations. Our framework identifies key components of stereotypes that are crucial in AI evaluation, including the target group, associated attribute, relationship characteristics, perceiving group, and relevant context. We also provide considerations and recommendations for its responsible use.
@@ 1 Introduction & Motivation
Recent years have seen unprecedented gains in generative AI models’ capabilities across modalities — language Anil et al. (2023); Achiam et al. (2023), image Rombach et al. (2022); Saharia et al. (2022), audio Kreuk et al. (2022); Borsos et al. (2023), and video Ho et al. (2022); Bar-Tal et al. (2024), while simultaneously gaining traction in diverse application domains and usage contexts across the globe Sengar et al. (2024); Raiaan et al. (2024). Along with these advancements, there are growing concerns that these models may reflect, propagate, and amplify societal stereotypes in their predictions and generations Garg et al. (2018a); Blodgett et al. (2020); Dev et al. (2022); Hovy and Prabhumoye (2021), and how they may lead to downstream harms Field et al. (2021); Shelby et al. (2023).
@@
A growing body of empirical work shows how NLP models reflect societal stereotypes about various groups — gender Bolukbasi et al. (2016), race Sap et al. (2019), nationality Jha et al. (2023), disability Hutchinson et al. (2020), to cite a few. Many of them also build datasets to enable large-scale evaluation of stereotypes in model predictions Nadeem et al. (2021); Jha et al. (2023); Bhutani et al. (2024). However, current research and resources lack a unified approach towards stereotypes in AI, hindering a comprehensive understanding of the problem space, thereby limiting effective and scalable interventions. First, they fail to capitalize on the underlying common mechanisms that may be contributing to stereotypes in society, data, and models. Additionally, it makes it harder to envision a unified way to tackle and prioritize downstream sociotechnical harms; instead potentially leading to unintended consequences, like new stereotypes emerging when others are mitigated. Another gap stems from adopting simplistic representations of stereotypes for expediency in evaluations, e.g., (identity, attribute) pairs overlook core aspects such as how stereotypes tie to specific time and place, which social groups hold certain stereotypes, and what connotations they imply.
@@
Finally, there are different methodologies to source stereotype data — e.g., annotator-driven collection Nadeem et al. (2021), LLM-enabled collection Jha et al. (2023), and community centered collection Dev et al. (2023a), each having unique strengths in terms of scalability, coverage, and reliability. However, we currently do not have an effective approach to determine which of these methods are appropriate in which contexts, what their relative merits (and demerits) are, and how to use these approaches in ways that lean on their strengths and complementarities. Having a unified framework will enable effective intervention, prioritization in high-stake environments, shared knowledge and methods across various efforts to collect data and intervene on models, predictions, and evaluations. Such a framework will also reveal aspects of this problem space that we still have large gaps to fill.
@@
In order to address these needs, we build off of social scientific theories on stereotypes as well as existing research on evaluating language technologies for stereotypes, and propose a unified comprehensive framework to operationalize stereotype evaluations. Our framework identifies various high level components such as the target group, the attribute associated with the group, the characteristics of their association, the perceiving group, as well as the context within which these stereotypes are prevalent. We also outline a set of recommendations for how to factor in responsibility considerations while using this framework.
@@ 2 Background
Social scientists have dedicated substantial research to the study of stereotypes, recognizing their intricate and multifaceted nature Macrae et al. (1996); Schneider (2005). This exploration has led to the development of various frameworks over time, aiming to unravel the complexities of how stereotypes originate, function, and influence both individuals and society as a whole Hilton and Von Hippel (1996). Early work predominantly viewed stereotypes as inaccurate generalizations about groups, stemming from limited or biased information Allport et al. (1954). They are also seen as cognitive shortcuts that help individuals simplify and categorize the social world, although this simplification could lead to errors and biases Dovidio et al. (2010). While these cognitive processes can be efficient, the connection between stereotypes (cognitive bias), prejudice (attitude bias), and discrimination (behavioral bias) was recognized early on, pointing to stereotypes as the motivation for negative attitudes and behaviors towards out-groups Macrae and Bodenhausen (2000).
@@
Various theories have been developed that focus on diverse aspects of stereotypes. Social identity theory emphasizes the role of group membership in shaping self-concept and inter-group relations, suggesting that stereotypes can serve to enhance one’s own group identity Tajfel et al. (1979). Social learning theory, on the other hand, focuses on stereotypes being learned through observation and socialization, often from parents, peers, and media Bandura and Walters (1977). System justification theory examines how stereotypes can be used to justify existing social hierarchies, even by members of disadvantaged groups Jost and Banaji (1994). Intersectionality theory further emphasizes the interconnected nature of social identities and how multiple stereotypes can intersect to create unique experiences of discrimination Crenshaw (2013).
@@
These theoretical perspectives have guided the development of various frameworks for analyzing stereotypes. Primarily shaped by social psychologists, these frameworks are widely used in other fields to model group dynamics and interactions. One of the prominent such frameworks is the Stereotype Content Model (SCM), which posits that stereotypes vary along two dimensions: warmth and competence, resulting in different emotional and behavioral responses towards groups Cuddy et al. (2007); Fiske et al. (2018). By extending the SCM, the dual perspectives model (Abele et al., 2016) added Morality and Sociability axes to the Warmth, and Ability and Assertiveness axes to the Competence dimension. Agency-Beliefs-Communion (ABC; Koch et al., 2016) model further added Status to the Competence dimension, and Belief as a dimension; “one end of Beliefs represents all religious, conservative, and other traditional groups; at the other end are progressives, artists, scientists, and LGBTQ groups.” Nicolas et al. (2022) relied on natural language processing approaches to both validate the SCM’s dimensions as well as discovering dimensions not commonly covered by SCM, such as Health, and Appearance.
@@
Some of these frameworks are increasingly being explored in NLP research. For instance, SCM has been applied to understand annotator biases Davani et al. (2023) and debiasing word embeddings Ungless et al. (2022); Omrani et al. (2023). Fraser et al. (2022) present a computational method to apply SCM to textual data and demonstrated that stereotypes in textual resources compare favourably with survey-based studies in the psychological literature. Fraser et al. (2024) used the ABC dimensions to evaluate and compare biases toward occupational groups across traditional survey-based data and various text sources. As NLP efforts increasingly grapple with the complexities of stereotypes in language, relying solely on social psychological frameworks of stereotypes can limit the scope of the analyses. These frameworks often prioritize dimensions like warmth and competence, potentially overlooking crucial aspects such as social dynamics, socio-historical context, and linguistic valence, which are also essential for a comprehensive understanding of stereotypes in language technologies.
@@ 3.1 Stereotype Detection and Evaluation
A significant number of responsible AI and NLP evaluations are concerned with various concepts that are inherently intertwined with stereotypes. For instance, bias measurement in co-reference resolution tasks often relies on gender-based occupation stereotypes Zhao et al. (2018); Rudinger et al. (2018), hate speech detection can hinge on societal stereotypes Chiril et al. (2021), offensive text can be comprised of stereotypes Jeong et al. (2022), sentiments that are disparately associated with different target groups stem from stereotypical perceptions about them Kiritchenko and Mohammad (2018), and more. However, the stereotype resources that these evaluations depend on, are limited in which groups they represent. While substantial work has focused on gender and racial stereotypes, they are also mostly constrained by binary gender constructs Dev et al. (2021) and Western racial histories Sambasivan et al. (2021). Other identity axes such as disability status, or socio-economic conditions are not as well represented. These resources are also strife with Western gaze wherein a majority of the resources are collected in the West (or even specifically North America), with data and annotators both representing Western viewpoints.
Based on keyword-based querying of the ACL anthology, [^1] https://aclanthology.org/ we note that 4140 papers mention stereotypes, their detection, resources, and evaluation. Of these, 54.1% mention gender-based stereotypes, 25.8% mention racial stereotypes, and only 16.4% mention region- and nationality-based stereotypes, and an even smaller fraction mention other identities such as age, disability, and profession. Some papers categorize stereotypes as positive or negative, often discussing the associated sentiment rather than the effect it can have downstream, or the specific marginalization the target groups experience Blodgett et al. (2021). For example, ‘women are polite’ can arguably be considered positive because of the sentiment associated with politeness, but the stereotype can have other implicit harms Cheng et al. (2023) related to the history of expectations of politeness and servitude from women Garg et al. (2018b), something that can negatively influence applications such as job recommendations based on gender.
@@ 3.2 Stereotype Resource Creation
Evaluating how stereotypes impact NLP model outputs requires societal data that capture such stereotypes. In this section, we will discuss different approaches used to build such datasets employed in NLP research.
@@
Historically, social psychology studies have provided a rich source of societal stereotypes that have been utilized to develop both resources and evaluation strategies for AI models Caliskan et al. (2017). These studies can contribute societal grounding regarding how a stereotype is perceived Fiske (1991); Kite et al. (2022), as well as provide extensive examples of prevalent stereotypes about different groups of people Borude (1966) that have been used in NLP evaluations Bhatt et al. (2022), and even lead to fine-tuning existing stereotype conntent models to LLM setting (Nicolas and Caliskan, 2024b, a).
@@
Crowdsourced data, while exceptionally valuable, is often restricted in its media form (primarily text), representation (who participates in crowdsourcing), and time (reflecting a specific moment). Researchers, therefore, turned to “big data” resources (e.g., social networks, and web crawls) which offer a broader range of content, perspectives, and temporal data. Existing media content, whether text, images, or videos, is shown reflect the stereotypes present in the society. Wikipedia, for instance, documents the origins of some well-known stereotypes and describes their provenance. News articles and social media can (un)intentionally propagate stereotypes as expressed by their authors. A popular approach for collecting such stereotypes is to crawl resources and capture co-occurrences of identity terms and attribute words (Sap et al., 2020; Bhatt et al., 2022; Bourgeade et al., 2023).
@@
While crowdsourcing and social media based curation increase the scale of stereotype resources, they are still limited in coverage of identities and range of associated stereotypes. More recent approaches has looked into leveraging large language models to expand coverage of stereotypes in a rapidly scalable manner and create a resource with broader coverage. When coupled with human annotations, these approached provide validated resources that even significantly overcome selection bias of data creators  Jha et al. (2023); Bhutani et al. (2024). While this expands the state of stereotype resources across identity axes, languages, and cultures, such an approach holds only when models are exposed (via their training data) to such social information in specific languages and about particular identity groups; thus leaving gaps in coverage across the world and many marginalized groups who are not well-represented in the online discourse.
@@
These approaches often offer complementary strengths and weaknesses Dev et al. (2023b). E.g., social psychological studies and community sourced studies tend to generate relatively smaller resources, but they bring forth richer and nuanced perspectives such as the perceiver group, and the extent of marginalization of the target group, while filling gaps about communities that are underrepresented in existing resources.
@@ 3.3 Gaps in Current Approaches
While the variety of approaches for collecting stereotypes do overlap and address some gaps (e.g., scalability and coverage), significant limitations persists across many of the mentioned approaches.
@@
Stereotypes are not static but rather temporally variable. They are influenced by how terms get reclaimed and change in meaning, historical events that lead to a shift in sentiment towards groups of people, and more (e.g., Garg et al. (2018b)). Yet, most resources capture stereotypes as a snapshot without capturing their evolving nature. For a resource to be operationalizable in bias mitigation or data and model evaluations, temporal grounding is critical. This helps resolve questions regarding factuality (e.g., French kings in 1600s being White is factual and not stereotypical) and misinformation (current Pope is not female, or Asians being associated with COVID 19 post the pandemic Lin et al. (2022)), identification of offensive slurs or pejorative terms (e.g., the word protestant was derogatory in 1500s but is simply a descriptor of religious identity now) and prevalent discriminatory practices (e.g. fraction of women who could vote in the United States before and after the women’s suffrage movement Garg et al. (2018b)).
Silowed Stereotype Evaluations:
@@
As discussed by Blodgett et al. (2021) in a thorough assessment of a number NLP measurements of stereotypes, benchmarks do not always rely on solid conceptualizations of stereotypes. Definitions of stereotype often lack critical components such as power dynamics, and consistency in defining social categories. Moreover, even thorough considerations during conceptualization are not guaranteed to be accurately reflected into operationalization. While these gaps are often hard to completely eliminate, it is important to articulate them to further attention on more effective operationalizations.
@@
While stereotypes are born as interactions between social groups, one being the group that is perceiving and one the group that is being perceived, most frameworks and benchmarks do not consider the perceiver group and solely focus on the target group. Notably, Jha et al. (2023) point out that individuals in different geographic regions are familiar with different non-overlapping stereotypes about the same identities. While computational work on stereotypes have expanded the participants pool through crowdsourcing — although the intention for this is often to reduce the cost and time, and not to diversify the sample, they still do not take the crowdworkers background information into account in how these resources are used.
@@
Not every over-represented association is a stereotype. Stereotypes require societal grounding for identification of harms caused Bhatt et al. (2022). Large-scale model evaluations for stereotypical or ‘biased’ behavior without contextual grounding merely calibrates model tendencies. A common example is racial bias and specifically anti-African-American stereotypes that are prevalent in the United States and rooted in colonial history, but are not as prevalent in South Asia where skin color does not correlate with race or nationality. Grounding a stereotype with what specific socio-cultural settings it is common in, helps build better evaluation paradigms and generative AI systems Sambasivan et al. (2021).

### v3

@@ Abstract
Societal stereotypes are at the center of a myriad of responsible AI interventions targeted at reducing the generation and propagation of potentially harmful outcomes. While these efforts are much needed, they tend to be fragmented and often address different parts of the issue without adopting a unified or holistic approach to social stereotypes and how they impact various parts of the machine learning pipeline. As a result, current interventions fail to capitalize on the underlying mechanisms that are common across different types of stereotypes, and to anchor on particular aspects that are relevant in certain cases. In this paper, we draw on social psychological research and build on NLP data and methods, to propose a unified framework to operationalize stereotypes in generative AI evaluations. Our framework identifies key components of stereotypes that are crucial in AI evaluation, including the target group, associated attribute, relationship characteristics, perceiving group, and context. We also provide considerations and recommendations for its responsible use.
@@ 1 Introduction & Motivation
Recent years have seen unprecedented gains in generative AI models’ capabilities across modalities—language Anil et al. (2023); Achiam et al. (2023), image Rombach et al. (2022); Saharia et al. (2022), audio Kreuk et al. (2022); Borsos et al. (2023), and video Ho et al. (2022); Bar-Tal et al. (2024)—while simultaneously gaining traction in diverse application domains and usage contexts across the globe Sengar et al. (2024); Raiaan et al. (2024). Along with these advancements, there are growing concerns that these models may reflect, propagate, and amplify societal stereotypes in their predictions and generations Garg et al. (2018a); Blodgett et al. (2020); Dev et al. (2022); Hovy and Prabhumoye (2021), potentially leading to downstream harms Field et al. (2021); Shelby et al. (2023).
@@
A growing body of empirical work shows how NLP models reflect societal stereotypes about various groups—including gender Bolukbasi et al. (2016), race Sap et al. (2019), nationality Jha et al. (2023), and disability Hutchinson et al. (2020) to cite a few. Many efforts also build datasets to enable large-scale evaluation of stereotypes in model predictions Nadeem et al. (2021); Jha et al. (2023); Bhutani et al. (2024). However, current research and resources lack a unified approach toward stereotypes in AI, hindering a comprehensive understanding of the problem space and, thereby, limiting effective and scalable interventions. First, they fail to capitalize on the underlying common mechanisms that may be contributing to stereotypes in society, data, and models. Consequently, it makes it harder to envision a unified way to tackle and prioritize downstream sociotechnical harms; which could instead lead to unintended consequences, like new stereotypes emerging when others are mitigated. Another gap stems from adopting simplistic representations of stereotypes for expediency in evaluations, e.g., (identity, attribute) pairs overlook core aspects such as how stereotypes tie to specific time and place, which social groups hold certain stereotypes, and what connotations they imply.
@@
Finally, there are different methodologies to source stereotype data—e.g., annotator-driven collection Nadeem et al. (2021), LLM-enabled collection Jha et al. (2023), and community-centered collection Dev et al. (2023a)—each having unique strengths in terms of scalability, coverage, and reliability. However, we currently do not have an effective approach to determine which of these methods are appropriate in which contexts, what their relative merits (and demerits) are, and how to use these approaches in ways that lean on their strengths and complementarities. Having a unified framework will enable effective intervention, prioritization in high-stake environments, shared knowledge and methods across various efforts to collect data and intervene on models, predictions, and evaluations. Such a framework will also reveal aspects of this problem space that we still have large gaps to fill.
@@
In order to address these needs, we build off of social scientific theories on stereotypes as well as existing research on evaluating language technologies for stereotypes, and propose a unified, comprehensive framework to operationalize stereotype evaluations. Our framework identifies various high level components such as the target group, the attribute associated with the group, the characteristics of their association, the perceiving group, as well as the context within which these stereotypes are prevalent. We also outline a set of recommendations for how to factor in responsibility considerations while using this framework.
@@ 2 Background
Social scientists have dedicated substantial research to the study of stereotypes, recognizing their intricate and multifaceted nature Macrae et al. (1996); Schneider (2005). This exploration has led to the development of various frameworks over time, aiming to unravel the complexities of how stereotypes originate, function, and influence both individuals and society as a whole Hilton and Von Hippel (1996). Early work predominantly viewed stereotypes as inaccurate generalizations about groups, stemming from limited or biased information Allport et al. (1954). Stereotypes are also seen as cognitive shortcuts that help individuals simplify and categorize the social world, although this simplification could lead to errors and biases Dovidio et al. (2010). While these cognitive processes can be efficient, the connection between stereotypes (cognitive bias), prejudice (attitude bias), and discrimination (behavioral bias) was recognized early on, pointing to stereotypes as the motivation for negative attitudes and behaviors toward out-groups Macrae and Bodenhausen (2000).
@@
Various theories have been developed that focus on diverse aspects of stereotypes. The Social identity theory emphasizes the role of group membership in shaping self-concept and inter-group relations, suggesting that stereotypes can serve to enhance one’s own group identity Tajfel et al. (1979). The Social learning theory, on the other hand, focuses on stereotypes being learned through observation and socialization, often from parents, peers, and media Bandura and Walters (1977). The System justification theory examines how stereotypes can be used to justify existing social hierarchies, even by members of disadvantaged groups Jost and Banaji (1994). Finally, Intersectionality theory further emphasizes the interconnected nature of social identities and how multiple stereotypes can intersect to create unique experiences of discrimination Crenshaw (2013).
@@
These theoretical perspectives have guided the development of various frameworks for analyzing stereotypes. Primarily shaped by social psychologists, these frameworks are widely used in other fields to model group dynamics and interactions. One of the prominent such frameworks is the Stereotype Content Model (SCM), which posits that stereotypes vary along two dimensions: Warmth and Competence, resulting in different emotional and behavioral responses toward groups Cuddy et al. (2007); Fiske et al. (2018). By extending the SCM, the dual perspectives model (Abele et al., 2016) added Morality and Sociability axes to the Warmth, and Ability and Assertiveness axes to the Competence dimension. Agency-Beliefs-Communion (ABC; Koch et al., 2016) model further added Status to the Competence dimension and Belief as a dimension; specifically, “one end of Beliefs represents all religious, conservative, and other traditional groups; at the other end are progressives, artists, scientists, and LGBTQ groups.” Nicolas et al. (2022) relied on natural language processing approaches to both validate the SCM’s dimensions as well as discovering dimensions not commonly covered by SCM, such as Health and Appearance.
@@
Some of these frameworks are increasingly being explored in NLP research. For instance, SCM has been applied to understand annotator biases Davani et al. (2023) and debiasing word embeddings Ungless et al. (2022); Omrani et al. (2023). Fraser et al. (2022) present a computational method to apply SCM to textual data and demonstrated that stereotypes in textual resources compare favorably with survey-based studies in the psychological literature. Fraser et al. (2024) used the ABC dimensions to evaluate and compare biases toward occupational groups across traditional survey-based data and various text sources. As NLP efforts increasingly grapple with the complexities of stereotypes in language, relying solely on social psychological frameworks of stereotypes can limit the scope of the analyses. These frameworks often prioritize dimensions like warmth and competence, potentially overlooking crucial aspects such as social dynamics, socio-historical context, and linguistic valence, which are also essential for a comprehensive understanding of stereotypes in language technologies.
@@ 3.1 Stereotype Detection and Evaluation
A significant number of responsible AI and NLP evaluations are concerned with various concepts that are inherently intertwined with stereotypes. For instance, bias measurement in co-reference resolution tasks often relies on gender-based occupation stereotypes Zhao et al. (2018); Rudinger et al. (2018); hate speech detection can hinge on societal stereotypes Chiril et al. (2021); offensive text can be comprised of stereotypes Jeong et al. (2022); sentiments that are disparately associated with different target groups stem from stereotypical perceptions about them Kiritchenko and Mohammad (2018); and more. However, the stereotype resources that these evaluations depend on, are limited in which groups they represent. While substantial work has focused on gender and racial stereotypes, they are also mostly constrained by binary gender constructs Dev et al. (2021) and Western racial histories Sambasivan et al. (2021). Other identity axes such as disability status or socio-economic conditions are not as well represented. These resources are also rife with Western gaze wherein a majority of the resources are collected in the West (or even specifically North America), with data and annotators both representing Western viewpoints.
Based on keyword-based querying of the ACLanthology, [^1] https://aclanthology.org/ we note that 4140 papers mention stereotypes, their detection, resources, and evaluation. Of these, 54.1% mention gender-based stereotypes, 25.8% mention racial stereotypes, and only 16.4% mention region- and nationality-based stereotypes, and an even smaller fraction mention other identities such as age, disability, and profession. Some papers categorize stereotypes as positive or negative, often discussing the associated sentiment rather than the effect it can have downstream or the specific marginalization the target groups experience Blodgett et al. (2021). For example, “women are polite” can arguably be considered positive because of the sentiment associated with politeness, but the stereotype can have other implicit harms Cheng et al. (2023) related to the history of expectations of politeness and servitude from women Garg et al. (2018b), something that can negatively influence applications such as job recommendations based on gender.
@@ 3.2 Stereotype Resource Creation
Evaluating how stereotypes impact NLP model outputs requires societal data that capture such stereotypes. In this section, we discuss different approaches used to build such datasets employed in NLP research.
@@
Historically, social psychology studies have provided a rich source of societal stereotypes that have been utilized to develop both resources and evaluation strategies for AI models Caliskan et al. (2017). These studies can contribute societal grounding regarding how a stereotype is perceived Fiske (1991); Kite et al. (2022), as well as provide extensive examples of prevalent stereotypes about different groups Borude (1966) that have been used in NLP evaluations Bhatt et al. (2022), and even lead to fine-tuning existing stereotype content models to LLM setting (Nicolas and Caliskan, 2024b, a).
@@
Crowdsourced data, while exceptionally valuable, is often restricted in its media form (primarily text), representation (who participates in crowdsourcing), and time (reflecting a specific moment). Researchers, therefore, turned to “big data” resources (e.g., social networks, and web crawls) which offer a broader range of content, perspectives, and temporal data. Existing media content, whether text, images, or videos, is shown to reflect the stereotypes present in the society. Wikipedia, for instance, documents the origins of some well-known stereotypes and describes their provenance. News articles and social media can propagate stereotypes as expressed by their authors. A popular approach for collecting such stereotypes is to crawl resources and capture co-occurrences of identity terms and attributes (Sap et al., 2020; Bhatt et al., 2022; Bourgeade et al., 2023).
@@
While crowdsourcing and social media based curation increase the scale of stereotype resources, they are still limited in coverage of identities and range of associated stereotypes. More recent approaches have looked into leveraging large language models to expand coverage of stereotypes in a rapidly scalable manner and create a resource with broader coverage. When coupled with human annotations, these approaches provide validated resources that even significantly overcome selection bias of data creators  Jha et al. (2023); Bhutani et al. (2024). While this expands the state of stereotype resources across identity axes, languages, and cultures, such an approach holds only when models are exposed (via their training data) to such social information in specific languages and about particular identity groups; thus leaving gaps in coverage across the world and many marginalized groups who are not well-represented in the online discourse.
@@
These approaches often offer complementary strengths and weaknesses Dev et al. (2023b). For instance, social psychological studies and community sourced studies tend to generate relatively smaller resources, but they bring forth richer and nuanced perspectives such as the perceiver group, and the extent of marginalization of the target group, while filling gaps about communities that are underrepresented in existing resources.
@@ 3.3 Gaps in Current Approaches
While the variety of approaches for collecting stereotypes do overlap and address some gaps (e.g., scalability and coverage), significant limitations persist across many of the mentioned approaches.
@@
Stereotypes are not static but rather temporally variable. They are influenced by how terms get reclaimed and change in meaning, historical events that lead to a shift in sentiment toward groups of people, and more (e.g., Garg et al. (2018b)). Yet, most resources capture stereotypes as a snapshot without capturing their evolving nature. For a resource to be operationalizable in bias mitigation or data and model evaluations, temporal grounding is critical. This helps resolve questions regarding factuality (e.g., French kings in 1600s being White is factual and not stereotypical) and misinformation (current Pope is not female, or Asians being associated with COVID 19 post the pandemic Lin et al. (2022)), identification of offensive slurs or pejorative terms (e.g., the word Protestant was derogatory in 1500s but is simply a descriptor of religious identity now) and prevalent discriminatory practices (e.g., fraction of women who could vote in the United States before and after the women’s suffrage movement Garg et al. (2018b)).
Siloed Stereotype Evaluations:
@@
As discussed by Blodgett et al. (2021) in a thorough assessment of a number NLP measurements of stereotypes, benchmarks do not always rely on solid conceptualizations of stereotypes. Definitions of stereotype often lack critical components such as power dynamics and consistency in defining social categories. Moreover, even thorough considerations during conceptualization are not guaranteed to be accurately reflected into operationalization. While these gaps are often hard to completely eliminate, it is important to articulate them to further focus on more effective operationalizations.
@@
While stereotypes are born as interactions between social groups, one being the group that is perceiving and one the group that is being perceived, most frameworks and benchmarks do not consider the perceiver group and solely focus on the target group. Notably, Jha et al. (2023) point out that individuals in different geographic regions are familiar with different non-overlapping stereotypes about the same identities. While computational work on stereotypes have expanded the participant pool through crowdsourcing—although the intention for this is often to reduce the cost and time, and not to diversify the sample, they still do not take the crowdworkers background information into account in how these resources are used.
@@
Not every over-represented association is a stereotype. Stereotypes require societal grounding for identification of harms caused Bhatt et al. (2022); Zhou et al. (2023). Large-scale model evaluations for stereotypical or “biased” behavior without contextual grounding merely calibrate model tendencies. A common example is racial bias and specifically anti-African-American stereotypes that are prevalent in the United States and rooted in colonial history, but are not similarly prevalent in South Asia where skin color does not correlate with race or nationality. Grounding a stereotype with what specific socio-cultural settings it is common in, helps build better evaluation paradigms and generative AI systems Sambasivan et al. (2021).